The collector current of a transistor used as an amplifier changes from 6MA to 18MA as the base current change from 5µA to 25µA. What is the current amplification factor of a transistor?
1
Expert's answer
2020-06-26T14:37:21-0400

By definition, the current amplification factor of a transistor is:


"\beta = \dfrac{\Delta I_{coll}}{\Delta I_{base}}"

where "\Delta I_{coll}" is change in the collector current when the base current changes by "\Delta I_{base}".

Substituting values, get:


"\beta = \dfrac{(18-6)\cdot 10^{-3}A}{(25-5)\cdot 10^{-6}A} = 600"

Answer. The current amplification factor of a transistor is 600.